OpenStudy (shubhamsrg):

summation (1 /(n^3 -n) ) from n=2 to infinity

OpenStudy (shubhamsrg):

I have been able to simplify it, 1/2(n+1) + 1/2(n-1) -1/n but stuck after this

OpenStudy (shubhamsrg):

I already know the answer, don;t know how to reach there. I don;t see a pattern when I plug in values of n

OpenStudy (anonymous):

these is the harmonis series \[H=\sum_{k=1}^{\infty}\frac{ 1 }{ k }\]

OpenStudy (anonymous):

How about splitting it like this\[\frac{A}{n(n-1)}+\frac{B}{n(n+1)}\]

OpenStudy (phi):

replace sum n=2 to inf with n=3 to inf for 1/2(n+1) replace sum n=2 to inf with n=1 to inf for 1/2(n-1) then "pull out" any terms less than n=3 to inf. If that makes sense.....

OpenStudy (phi):

\[\sum_{2}^{\infty}\frac{1}{2(n+1)}= \frac{1}{2}\sum_{3}^{\infty}\frac{1}{n}\]

OpenStudy (shubhamsrg):

oh you mean 1/n = 1/2n + 1/2n

OpenStudy (shubhamsrg):

I see we can represent the whole series as sigma(1/n) but won't we get stuck after that ?

OpenStudy (phi):

\[\sum_{2}^{\infty}\frac{1}{2(n-1)}= \frac{1}{2}\sum_{1}^{\infty}\frac{1}{n}=\frac{1}{2}+\frac{1}{4}+ \frac{1}{2}\sum_{3}^{\infty}\frac{1}{n}\]

OpenStudy (shubhamsrg):

@mukushla sire thats what you meant ? split 1/n in 1/2n + 1/2n ?

OpenStudy (shubhamsrg):

@phi but we do not know the finite sum of that series ?

OpenStudy (phi):

\[-\sum_{2}^{\infty}\frac{1}{n}= \frac{1}{2}-\sum_{3}^{\infty}\frac{1}{n}\]

OpenStudy (phi):

now combine the summations \[\frac{1}{2}+\frac{1}{4}-\frac{1}{2}+\frac{1}{2}\sum_{3}^{\infty}\frac{1}{n}+\frac{1}{2}\sum_{3}^{\infty}\frac{1}{n}-\sum_{3}^{\infty}\frac{1}{n}\]

OpenStudy (phi):

notice that the summations add to zero.

OpenStudy (phi):

you are left with 1/4

OpenStudy (shubhamsrg):

oh yes, nice ! thank you @phi

OpenStudy (shubhamsrg):

how did that strike you ?

OpenStudy (anonymous):

\[\huge \sum_{n=1}^{\infty} (\frac{ 1 }{ 2n(n+1) }+\frac{-1 }{2n(n-1)})\] from @mukushla there seem to be a cancellıng of temrs

OpenStudy (shubhamsrg):

yes, both methods are brilliant! I understand both thank you @mukushla and @phi (:

OpenStudy (anonymous):

yw :)